OpenStudy (anonymous):

If X#Y=X-Y+2 then find the value of 2 # (3#5)? A) 6 B) 4 C) 2 D) 0

OpenStudy (anonymous):

Could someone pliz check if my answer is rite? (3#5)=3-5+2 5-5=0 2 # (0)=2-0+2 =4

OpenStudy (lgbasallote):

yup correct

OpenStudy (anonymous):

thanx :)

OpenStudy (lgbasallote):

<tips imaginary hat>

OpenStudy (anonymous):

which means?

OpenStudy (lgbasallote):

haha nothing

OpenStudy (anonymous):

oh yeah i knew that